# Learning-Augmented Skip Lists

### v1

@@ Abstract
We study the integration of machine learning advice into the design of skip lists to improve upon traditional data structure design. Given access to a possibly erroneous oracle that outputs estimated fractional frequencies for search queries on a set of items, we construct a skip list that provably provides the optimal expected search time, within nearly a factor of two. In fact, our learning-augmented skip list is still optimal up to a constant factor, even if the oracle is only accurate within a constant factor. We show that if the search queries follow the ubiquitous Zipfian distribution, then the expected search time for an item by our skip list is only a constant, independent of the total number $n$ of items, i.e., $\mathcal{O}(1)$, whereas a traditional skip list will have an expected search time of $\mathcal{O}(\log n)$. We also demonstrate robustness by showing that our data structure achieves an expected search time that is within a constant factor of an oblivious skip list construction even when the predictions are arbitrarily incorrect. Finally, we empirically show that our learning-augmented skip list outperforms traditional skip lists on both synthetic and real-world datasets.
@@ 1 Introduction
As efficient data management has become increasingly crucial, the integration of machine learning (ML) has significantly improved the design and performance of traditional algorithms for many big data applications. [${}^{+}$18] first showed that ML could be incorporated to create data structures that support faster look-up operations while also saving an order-of-magnitude of memory compared to optimized data structures oblivious to such ML heuristics. Subsequently, learning-augmented algorithms [MV20] have been shown to achieve provable worst-case guarantees beyond the limitations of oblivious algorithms for a wide range of settings. For example, ML predictions have been utilized to achieve more efficient data structures [Mit18, LLW22], algorithms with faster runtimes [${}^{+}$21, CSVZ22, DMVW23], mechanisms with better accuracy-privacy tradeoffs [KADV23], online algorithms with better performance than information-theoretic limits [PSK18, GP19, LLMV20, WLW20, WZ20, BMS20, AGP20, ${}^{+}$21, AGKP21, IKQP21, LV21, ACI22, AGKP22, APT22, ${}^{+}$22, KBTV22, ${}^{+}$22, SGM22, AGKK23, ${}^{+}$23, SLLA23], streaming algorithms with better accuracy-space tradeoffs [HIKV19, IVY19, ${}^{+}$20, CIW22, ${}^{+}$22, ${}^{+}$23], and polynomial-time algorithms beyond hardness-of-approximation limits, e.g., NP-hardness [${}^{+}$22, NCN23].
In this paper, we focus on the consolidation of ML advice to improve data structures for the fundamental problem of searching for ordered elements among a large dataset. For this purpose, tree-based structures stand out as a popular choice among other structures, particularly for their logarithmic average performance. However, these structures often need to be balanced for optimal performance, and thus their effectiveness is often closely tied to the order of element insertions. In contrast, skip lists, introduced by Pugh in 1990 [Pug90a], maintain balance probabilistically, offering a simpler implementation while delivering substantial speed enhancements [Pug90b]. Thus while binary search trees have been a long-standing choice for querying ordered elements, skip lists offer a simpler, more efficient, and in some cases necessary alternative.
Skip lists.
Skip lists are generally built iteratively in levels. The bottom levels of the skip list is an ordinary-linked list, in which the items of the dataset are organized in order. Each higher level serves to accelerate the search for the lower levels, by storing only a subset of the items in the lower levels, also as an ordered link list. Traditional skip lists are built by promoting each item in a level to a higher level randomly with a fixed probability $p\in(0,1)$.
@@
In many modern applications, skip lists are used because of their excellent search runtime and their space efficiency. Skip lists are often preferred over binary search trees due to their simplicity of implementation, their support for efficient range query, and their amenability to concurrent processes [SL00, LJ13], high efficiency for dynamic datasets [GZ08, PT10], network routing [HPJ03, ASS20], and real-time analytics [${}^{+}$20, ${}^{+}$23].
1.1 Skip Lists and Applications
Traditional skip lists treat each element equally when promoting the elements to higher levels. This balancing behavior facilitates good performance in expectation when a query to the skip list is equally likely to be any dataset element. On the other hand, this behavior may limit the performance of the data structure when the incoming queries are from an unbalanced probability distribution.
@@ 1.1 Skip Lists and Applications
In a Zipfian distribution, the frequency of an event $N(k;\alpha,N)$ is inversely proportional to its rank $k$, raised to the power of $\alpha$ (where $\alpha$ is a positive parameter), in a dataset of $N$ elements. The formula is given by:
$N(k;\alpha,N)=\frac{1/k^{\alpha}}{\sum_{n=1}^{N}(1/n^{\alpha})}$
The value of $\alpha$ determines the steepness of the distribution. A smaller $\alpha$ (close to 0) makes it more uniform, while a larger $\alpha$ increases skewness.
Zipfian distributions provide a simple means for understanding phenomena in various fields involving rank and frequency, ranging from linguistics to economics, and from urban studies to information technology. Indeed, they appear in many applications such as word frequencies in natural language [WW16, BHZ18], city populations [Gab99, VA15], biological cellular distributions [${}^{+}$23], income distribution [San15], etc.
Unfortunately, although Zipfian distributions are common in practice, their properties are not leveraged by traditional skip lists, which are oblivious to any information about the query distributions. To improve this performance bottleneck, we propose the augmentation of traditional skip lists with “learned” advice, which (possibly erroneously) informs the data structure in advance about some useful statistics on the incoming queries. Although we model the data structure as having oracle access to the advice, in practice, such advice can often easily be realized from machine learning heuristics trained for these useful statistics.
@@ 1.2 Our Contributions
We propose the incorporation of ML advice into the design of skip lists to improve upon traditional data structure design. In this setting, we assume the items that may appear either in the data set or the query set can be associated with an integer in $[n]:=\{1,\ldots,n\}$. We allow the algorithm access to a possibly erroneous oracle that, for each $i\in[n]$, outputs a quantity $p_{i}$, which should be interpreted as an estimation for the proportion of search queries that will be made to the data structure for the item $i$. Hence, for each $i\in[n]$, we assume that $p_{i}\in[0,1]$ and $p_{1}+\ldots+p_{n}=1$. Note that these constraints can be easily enforced upon the oracle as a pre-processing step prior to designing the skip list data structure. We also assume that the oracle is readily accessible so that there is no cost for each interaction with the oracle. Consequently, we assume the algorithm has access to the predicted frequency $p_{i}$ by the oracle for all $i\in[n]$. On the other hand, we view a sequence of queries as defining a probability distribution over the set of queries, so that $f_{i}$ is the true proportion of queries to item $i$, for each $i\in[n]$. Although $f_{i}$ is the ground truth, our algorithms only have access to $p_{i}$, which may or may not accurately capture $f_{i}$.
Consistency for accurate oracles.
@@
We describe a construction for a learning-augmented skip list, which gives expected search time at most $2C+2\sum_{i=1}^{n}f_{i}\cdot\min\left(\log\frac{1}{p_{i}},\log n\right)$, for some constant $C>0$. On the other hand, we show that any skip list construction requires an expected search time of at least the entropy $H(f)$ of the probability vector $f$. We recall that the entropy $H(f)$ is defined as $H(f)=\sum_{i=1}^{n}f_{i}\cdot\log\frac{1}{f_{i}}$.
@@
Thus, our results indicate that within nearly a factor of two, our learning-augmented skip list is optimal for any distribution of queries, provided that the oracle is perfectly accurate. Moreover, even if the oracle on each estimated probability $p_{i}$ is only accurate up to a constant factor, then our learning-augmented skip list is still optimal, up to a constant factor.
Implications to Zipfian distributions.
@@
We describe the implications of our results to queries that follow a Zipfian distribution; analogous results hold for other skewed distributions, e.g., the geometric distribution. It is known that if the $r$-th most common query/item has proportion $\frac{z}{r^{s}}$ for some $s>1$, then the entropy of the corresponding probability vector is a constant. Consequently, if the set of queries follows a Zipfian distribution and the oracle is approximately accurate within a constant factor, then the expected search time for an item by our skip list is only a constant, independent of the total number of items, i.e., $\mathcal{O}\left(1\right)$. By comparison, a traditional skip list will have expected search time $\mathcal{O}\left(\log n\right)$.
Robustness to erroneous oracles.
@@
So far, our discussions have centered around an oracle that either produces estimated probabilities $p_{i}$ such that $p_{i}=f_{i}$ or $p_{i}$ is within a constant factor of $f_{i}$. However, in some cases, the machine learning algorithm serving as the oracle can be completely wrong. In particular, a model that is trained on a dataset before a distribution change, e.g., seasonal trends or other temporal shifts, can produce wildly inaccurate predictions. We show that our skip list is robust to erroneous oracles. Specifically, we show that our algorithm achieves an expected search time that is within a constant factor of an oblivious skip list construction when the predictions are incorrect. Therefore, our data structure achieves both consistency, i.e., good algorithmic performance when the oracle is accurate, and robustness, i.e., standard algorithmic performance when the oracle is inaccurate.
Empirical evaluations.
@@
Finally, we analyze our learning-augmented skip list on both synthetic and real-world datasets. Firstly, we compare the performance of traditional skip lists with our learning-augmented skip lists on synthetically generated data following Zipfian distributions with various tail parameters. The dataset is created using four distinct $\alpha$ values ranging from 1.01 to 2, along with a uniform dataset. During the assessment, we query a specified number of $n$ items selectively chosen based on their frequency weights. Our results match our theory, showing that learning-augmented skip lists have faster query times, with an average speed-up factor ranging from 1.33 up to 7.76, depending on the different skewness parameters.
@@
1.3 Background on Skip Lists
The skip list data structure as discussed in Section 1.1 is a commonly adopted structure that that utilizes probabilistic attributes within the data to allow fast search within an ordered sequence of elements. Its efficiency stems from its layered structure, where each element is inserted into multiple sorted linked lists, with the top layers providing a fast path for traversing the list. Studies around skip lists have been focusing on techniques to improve their performance under specific use scenarios. [${}^{+}$19] proposed a layered framework structure to handle in-memory key-value indexing. They proposed an ensemble consisting of two layers, a cache-sensitive structure that maintains a few guard entries, and a Semi-Ordered Skip-List which supports high concurrency in insertions and offers fast lookup and range querying capabilities. This semi-ordered approach allows for efficient data handling, particularly beneficial for systems with heavy insertion workloads. [LJHY22] also improved upon traditional skip lists on persistent memory indexing applications. They introduced a hierarchical and partitioned skip list design, which reduces the height of the skip list and improves access locality. The skip list is organized into a hierarchy of component structures, which effectively enhances data access efficiency. [${}^{+}$19] also utilizes a neural network oracle for intelligent guard entry selection based on data and query distribution. This innovation ensures that the indexing structure adapts to the specific data characteristics, optimizing performance for various datasets. [${}^{+}$20] presents a novel skip list implementation to address the performance degradation issue with traditional skip lists in key-value stores (KVS) that use multi-version concurrency control (MVCC). [${}^{+}$20] reimagines a skip list by separating the per-key updates from the key indexing process, which reduces the overhead typically associated with MVCC in skip lists. Their evaluation demonstrates the efficiency improvements by up to 42% decrease in query latency in the context of RocksDB, making them more suitable for modern KVS environments.
To the best of our knowledge, this study is the first to extensively evaluate the performance of a learning augmented skip list with a novel implementation algorithm with a theoretical performance guarantee.
@@ 2.1 Algorithm
We first describe our learning-augmented skip list, which utilizes predictions $p_{i}$ for each item $i\in[n]$, from an oracle. Similar to a traditional skip list, the bottom level of our skip list is an ordinary-linked list that contains the sorted items of the dataset. As before, the purpose of each higher level is to accelerate the search for an item, but the process for promoting an item from a lower level to a higher level now utilizes the predictions. Whereas traditional skip lists promote each item in a level to a higher level randomly with a fixed probability $p\in(0,1)$, we automatically promote the item $i$ to a level $\ell$ if its predicted frequency $p_{i}$ satisfies $p_{i}\geq\frac{2^{\ell}-1}{n}$. Otherwise, we promote the item with probability $\frac{1}{2}$. This adaptation ensures that items with high predicted query frequencies will be promoted to higher levels of the skip list and thus be more likely to be found quickly. We give the full details in Algorithm 1.
It is worth noting that the implementation algorithm of the augmented skip lists has undergone several rounds of updates to eventually show its performance as shown herein. For example, in the first attempt of augmentation, the position with median probability is used as the pivot to promote elements to higher levels. This promotion scheme becomes problematic since the get median function could induce considerable time overhead for insertion operations. To resolve this issue, a separate threshold probability is set for each level so that only nodes with a probability higher than the corresponding threshold are promoted to the next level. This approach provides an alternative to the find_median function, which avoids the time overhead. Yet, this approach fails to cater to the diverse range of distributions and turns out to show unstable performance. To further resolve this issue, an ensemble promotion framework is proposed. The proposed framework first filters out the low-frequency elements, which remain on the bottom level. Then for each level of the promoted nodes, a threshold was determined dynamically according to the current level of the element, i.e., a threshold $\theta=f(level)$. A similar problem arises for the ensemble for extreme cases of distributions. For example, the ensemble may fail to level up the elements for a uniform distribution since all elements have the same probability, thus the dynamic threshold will fail to distinguish the elements to be promoted from those to remain. The strategy to solve this problem, which is proposed in our algorithm, is to add a guard for those elements that have been promoted from the bottom level. The guard ensures that those elements with higher probabilities will still have a chance of being promoted to higher levels even when their probability is less than the corresponding threshold. This updated ensemble framework is theoretically and experimentally guaranteed to perform at least as good as a vanilla skip list, and also outperform a vanilla skip list for specific types of distributions like a Zipfian distribution.

### v2

# Learning-Augmented Search Data Structures

@@ Abstract
We study the integration of machine learning advice to improve upon traditional data structure designed for efficient search queries. Although there has been recent effort in improving the performance of binary search trees using machine learning advice, e.g., Lin et. al. (ICML 2022), the resulting constructions nevertheless suffer from inherent weaknesses of binary search trees, such as complexity of maintaining balance across multiple updates and the inability to handle partially-ordered or high-dimensional datasets. For these reasons, we focus on skip lists and KD trees in this work. Given access to a possibly erroneous oracle that outputs estimated fractional frequencies for search queries on a set of items, we construct skip lists and KD trees that provably provides the optimal expected search time, within nearly a factor of two. In fact, our learning-augmented skip lists and KD trees are still optimal up to a constant factor, even if the oracle is only accurate within a constant factor. We also demonstrate robustness by showing that our data structures achieves an expected search time that is within a constant factor of an oblivious skip list/KD tree construction even when the predictions are arbitrarily incorrect. Finally, we empirically show that our learning-augmented search data structures outperforms their corresponding traditional analogs on both synthetic and real-world datasets.
@@ 1 Introduction
As efficient data management has become increasingly crucial, the integration of machine learning (ML) has significantly improved the design and performance of traditional algorithms for many big data applications. [KBC+18] first showed that ML could be incorporated to create data structures that support faster look-up operations while also saving an order-of-magnitude of memory compared to optimized data structures oblivious to such ML heuristics. Subsequently, learning-augmented algorithms [MV20] have been shown to achieve provable worst-case guarantees beyond the limitations of oblivious algorithms for a wide range of settings. For example, ML predictions have been utilized to achieve more efficient data structures [Mit18, LLW22], algorithms with faster runtimes [DIL+21, CSVZ22, DMVW23], mechanisms with better accuracy-privacy tradeoffs [KADV23], online algorithms with better performance than information-theoretic limits [PSK18, GP19, LLMV20, WLW20, WZ20, BMS20, AGP20, ACL+21, AGKP21, IKQP21, LV21, ACI22, AGKP22, APT22, GLS+22, KBTV22, JLL+22, SGM22, AGKK23, ACE+23, SLLA23], streaming algorithms with better accuracy-space tradeoffs [HIKV19, IVY19, JLL+20, CIW22, CEI+22, LLL+23], and polynomial-time algorithms beyond hardness-of-approximation limits, e.g., NP-hardness [EFS+22, NCN23, CLR+24].
In this paper, we focus on the consolidation of ML advice to improve data structures for the fundamental problem of searching for elements among a large dataset. For this purpose, tree-based structures stand out as a popular choice among other structures, particularly for their logarithmic average performance. However, these structures often have weaknesses for specific use cases that make them sub-optimal for various applications, which we now discuss.
Skip lists. One weakness of tree-based structures is that they need to be balanced for optimal performance, and thus their effectiveness is often closely tied to the order of element insertions. For example, the motivation of [LLW22] to study learning-augmented binary search trees noted that although previous results already characterized the statically optimal tree if the underlying distribution is known [Knu71, Meh77], these methods do not handle dynamic insertion operations. In contrast, skip lists, introduced by  [Pug90a], maintain balance probabilistically, offering a simpler implementation while delivering substantial speed enhancements [Pug90b]. Skip lists are generally built iteratively in levels. The bottom levels of the skip list is an ordinary-linked list, in which the items of the dataset are organized in order. Each higher level serves to accelerate the search for the lower levels, by storing only a subset of the items in the lower levels, also as an ordered link list. Traditional skip lists are built by promoting each item in a level to a higher level randomly with a fixed probability $p\in(0,1)$.
@@
In many modern applications, skip lists are used because of their excellent search runtime and their space efficiency. Skip lists are often preferred over binary search trees due to their simplicity of implementation, their support for efficient range query, and their amenability to concurrent processes [SL00, LJ13], high efficiency for dynamic datasets [GZ08, PT10], network routing [HPJ03, ASS20], and real-time analytics [BBB+20, ZCL+23]. Thus while binary search trees have been a long-standing choice for querying ordered elements, skip lists offer a simpler, more efficient, and in some cases, necessary alternative.
KD trees. Another weakness of tree-based data structures is that they generally require the data to obey an absolute ordering. However, in many cases, e.g., geometric applications or multidimensional data, the input points can only be partially ordered. Thus in 1975, KD trees, which stand for $k$-dimensional trees, were proposed as a more efficient alternative to binary search trees for searching in higher-dimensional spaces in procedures such as nearest neighbor search or ray tracing for applications in computational geometry or computer vision. A KD tree works by picking a data point and splitting along some spatial dimension to partition the space. This process is repeated until every data point is included in the tree, creating a hierarchical tree structure that enables quick access to specific data points or ranges within the dataset.
Skewed distributions. Traditional search data structures treat each element equally when promoting the elements to higher levels. This balancing behavior facilitates good performance in expectation when a query to the skip list is equally likely to be any dataset element. On the other hand, this behavior may limit the performance of the data structure when the incoming queries are from an unbalanced probability distribution.
@@
In a Zipfian distribution, the frequency of an event $N(k;\alpha,N)$ is inversely proportional to its rank $k$, raised to the power of $\alpha$ (where $\alpha$ is a positive parameter), in a dataset of $N$ elements. In particular, we have $N(k;\alpha,N)=\frac{1/k^{\alpha}}{\sum_{n=1}^{N}(1/n^{\alpha})}$. The value of $\alpha$ determines the steepness of the distribution so that a smaller $\alpha$ value, i.e., closer to 00, makes it more uniform, while a larger $\alpha$ increases skewness.
Zipfian distributions provide a simple means for understanding phenomena in various fields involving rank and frequency, ranging from linguistics to economics, and from urban studies to information technology. Indeed, they appear in many applications such as word frequencies in natural language [WW16, BHZ18], city populations [Gab99, VA15], biological cellular distributions [LVM+23], income distribution [San15], etc.
Unfortunately, although Zipfian distributions are common in practice, their properties are generally not leveraged by traditional search data structures, which are oblivious to any information about the query distributions. To improve this performance bottleneck, we propose the augmentation of traditional skip lists and KD trees with “learned” advice, which (possibly erroneously) informs the data structure in advance about some useful statistics on the incoming queries. Although we model the data structure as having oracle access to the advice, in practice, such advice can often easily be acquired from machine learning heuristics trained for these statistics.
@@ 1.1 Our Contributions
We propose the incorporation of ML advice into the design of skip lists and KD trees to improve upon traditional data structure design. For ease of discussion in this section, we assume the items that may appear either in the data set or the query set can be associated with an integer in $[N]:=\{1,\ldots,N\}$, which also in the case of high-dimensional data, may be associated with a $k$-dimensional point. We allow the algorithm access to a possibly erroneous oracle that, for each $i\in[N]$, outputs a quantity $p_{i}$, which should be interpreted as an estimation for the proportion of search queries that will be made to the data structure for the item $i$. Hence, for each $i\in[N]$, we assume that $p_{i}\in[0,1]$ and $p_{1}+\ldots+p_{n}=1$. Note that these constraints can be easily enforced upon the oracle as a pre-processing step prior to designing the skip list or KD tree data structure. We also assume that the oracle is readily accessible so that there is no cost for each interaction with the oracle. Consequently, we assume the algorithm has access to the predicted frequency $p_{i}$ by the oracle for all $i\in[N]$. On the other hand, we view a sequence of queries as defining a probability distribution over the set of queries, so that $f_{i}$ is the true proportion of queries to item $i$, for each $i\in[N]$. Although $f_{i}$ is the ground truth, our algorithms only have access to $p_{i}$, which may or may not accurately capture $f_{i}$.
@@
Consistency for accurate oracles. We introduce construction for a learning-augmented skip list and KD trees, which gives expected search time at most $2C+2\sum_{i=1}^{n}f_{i}\cdot\min\left(\log\frac{1}{p_{i}},\log n\right)$, for some constant $C>0$. On the other hand, we show that any skip list or KD tree construction requires an expected search time of at least the entropy $H(f)$ of the probability vector $f$. We recall that the entropy $H(f)$ is defined as $H(f)=\sum_{i=1}^{n}f_{i}\cdot\log\frac{1}{f_{i}}$.
@@
Thus, our results indicate that within nearly a factor of two, our learning-augmented search data structures are optimal for any distribution of queries, provided that the oracle is perfectly accurate. Moreover, even if the oracle on each estimated probability $p_{i}$ is only accurate up to a constant factor, then our learning-augmented search data structures are still optimal, up to a constant factor.
@@
Implications to Zipfian distributions. We describe the implications of our results to queries that follow a Zipfian distribution; analogous results hold for other skewed distributions, e.g., the geometric distribution. It is known that if the $r$-th most common query/item has proportion $\frac{z}{r^{s}}$ for some $s>1$, then the entropy of the corresponding probability vector is a constant. Consequently, if the set of queries follows a Zipfian distribution and the oracle is approximately accurate within a constant factor, then the expected search time for an item by our search data structures is only a constant, independent of the total number of items, i.e., $\mathcal{O}\left(1\right)$. By comparison, a traditional skip list or KD tree will have expected search time $\mathcal{O}\left(\log n\right)$.
@@
Robustness to erroneous oracles. So far, our discussions have centered around an oracle that either produces estimated probabilities $p_{i}$ such that $p_{i}=f_{i}$ or $p_{i}$ is within a constant factor of $f_{i}$. However, in some cases, the machine learning algorithm serving as the oracle can be completely wrong. In particular, a model that is trained on a dataset before a distribution change, e.g., seasonal trends or other temporal shifts, can produce wildly inaccurate predictions. We show that our search data structures are robust to erroneous oracles. Specifically, we show that our algorithms achieve an expected search time that is within a constant factor of an oblivious skip list or KD tree construction when the predictions are incorrect. Therefore, our data structure achieves both consistency, i.e., good algorithmic performance when the oracle is accurate, and robustness, i.e., standard algorithmic performance when the oracle is inaccurate.
@@
Empirical evaluations. Finally, we analyze our learning-augmented search data structures list on both synthetic and real-world datasets. Firstly, we compare the performance of traditional skip lists with our learning-augmented skip lists on synthetically generated data following Zipfian distributions with various tail parameters. The dataset is created using four distinct $\alpha$ values ranging from 1.01 to 2, along with a uniform dataset. During the assessment, we query a specified number of $n$ items selectively chosen based on their frequency weights. Our results match our theory, showing that learning-augmented skip lists have faster query times, with an average speed-up factor ranging from 1.33 up to 7.76, depending on the different skewness parameters.
@@
We similarly perform evaluations on our learning-augmented KD tree data structure. We evaluate our KD tree data structure on Zipfian distributions with various tail parameters, and provide a heatmap of average lookup times for elements. We find that for a large variety of Zipfian parameters, ourt method is able to provides large improvements over traditional KD trees. We perform a similar experiment under Zipfian distributions with added noise, and find our data structure still provides considerable improvements in query time.
We additionally evaluate our method on point cloud samples taken from a 3D model. We bin these samples in space, and create our learning-augmented KD tree on these binned samples. When querying this tree with new binned point samples, we find a decrease in average query time as compared to a traditional KD tree under the same conditions. In addition to this experiment, we provide results on real world datasets of n-grams and neuron activation, and similarly find improvements over traditional KD trees.
Concurrent and independent work. We mention that concurrent and independent of our work, [ZKH24] used similar techniques to achieve the same guarantees on the performance of learning-augmented skip lists that are robust to erroneous predictions. However, they do not show optimality for their learning-augmented skip lists and arguably perform less exhaustive empirical evaluations. They also do not consider KD trees at all, which forms a significant portion of our contribution, both theoretically and empirically.
Comparison to [LLW22]. Our work was largely inspired by [LLW22], who observed that classical literature characterizing statically optimal binary search trees [Knu71, Meh77] no longer apply in the dynamic setting, as elements arrive iteratively over time. Thus, they designed the construction of dynamic learning-augmented binary search trees (BSTs). Their analysis for the expected search time utilized the notion of pivots within their trees and thus were somewhat specialized to BSTs. Therefore, [LLW22] explicitly listed skip trees and advanced tree data structures as interesting open directions. Qualitatively, our results are similar to [LLW22], as are those of [ZKH24]. This is not quite altogether surprising because the main difference between these data structures is not necessarily the search time, but the either the ease of construction in the setting of skip lists, or the ability to handle multi-dimensional data in the setting of KD trees.
@@ 2 Learning-Augmented Skip Lists
We first describe our learning-augmented skip list, which utilizes predictions $p_{i}$ for each item $i\in[n]$, from an oracle. Similar to a traditional skip list, the bottom level of our skip list is an ordinary-linked list that contains the sorted items of the dataset. As before, the purpose of each higher level is to accelerate the search for an item, but the process for promoting an item from a lower level to a higher level now utilizes the predictions. Whereas traditional skip lists promote each item in a level to a higher level randomly with a fixed probability $p\in(0,1)$, we automatically promote the item $i$ to a level $\ell$ if its predicted frequency $p_{i}$ satisfies $p_{i}\geq\frac{2^{\ell}-1}{n}$. Otherwise, we promote the item with probability $\frac{1}{2}$. This adaptation ensures that items with high predicted query frequencies will be promoted to higher levels of the skip list and thus be more likely to be found quickly.
It is worth addressing a number of other natural approaches and their shortcomings. For example, one natural approach would be to use the “median” frequency across the items as a threshold to promote elements to higher levels. However, this promotion scheme is not ideal because computing the median frequency at each time would either require an additional data structure for fast update time or increase the insertion time. A potential approach to resolve this issue would be to use a separate threshold probability is set for each level so that only nodes with a probability higher than the corresponding threshold are promoted to the next level. However, this approach seems to result in an unnecessarily large number of created levels if some item appears with small probability, e.g., $\frac{1}{2^{n}}$. We can thus first filters out the low-frequency elements and place them remain on the bottom level of the skip list and then proceed with using a separate threshold probability for each level. Unfortunately, this approach utterly fails to even match the search time performance of oblivious skip lists when the distribution is uniform, because all items will be in the same level, resulting in an expected search time of $\Omega(n)$. Hence, we ensure that each element still has a chance of being promoted to higher levels even when their probability is less than the corresponding threshold.
We again emphasize that due to the dynamic nature of the updates, existing results on statically optimal binary search trees [Knu71, Meh77] do not apply, as observed by [LLW22]. We give the full details in Algorithm 1. For the sake of presentation, we focus on the setting where the queries are made to items in the dataset. However, we remark that our results generalize to the setting where queries can be made on the search space rather than the items in the dataset, provided the oracle is also appropriately adjusted to estimate the query distribution, using the approach we describe in Section 3.